OpenStudy (anonymous):

Find a Parametrization for the curve: The line through the points 3, 3 and 5, 1. Would the answer be: y = 8t+3 & x=4t-3

OpenStudy (tkhunny):

Do you mean "Would AN answer be...?" There is more than one parameterization. To get y = 0, your formulation requires t = 0. t = 0 gives x = -3 You seem to have missed the point! Try y = mt + 3 and x = nt + 3 Now, for t = 0, we hit (3,3) How do we choose n and m so as to find (5,1)?

OpenStudy (helder_edwin):

u have to points of the line u=(3,3) and v=(5,1) so the direction of the line is u-v=(3,3)-(5,1)=(-2,2) so u can write the line as \[ \large \{(3,3)+t(-2,2):t\in\mathbb{R}\} \] this means that \[ \large x=3-2t\qquad y=3+2t \]

OpenStudy (anonymous):

Yea I ment is my answer a possible answer

OpenStudy (helder_edwin):

u can check if your answer is right by solving \[ \large 3=8t+3\qquad 3=4t-3 \] u should get the same t from both equations.

OpenStudy (anonymous):

So it is right?